OpenStudy (anonymous):

It took John 12 hours riding his bike to make the round trip to his uncle's. If he averaged 20 mph out and 30 mph back, how long did he travel each way? (Round answer to nearest tenth.)

OpenStudy (anonymous):

guess we need distance is rate times time for this one time goes is \(\frac{d}{20}\) and time returning is \(\frac{d}{30}\)for a total time of \[\frac{d}{20}+\frac{d}{30}\]

OpenStudy (anonymous):

since you are told the total time is \(12\) sovlve \[\frac{d}{20}+\frac{d}{30}=12\]